OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3.

OpenStudy (anonymous):

y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

OpenStudy (luigi0210):

Well you have the slope, and a pair of coordinate points, just plug them into the point-slope formula \(y-y_{1}=m(x-x_{1})\)

OpenStudy (luigi0210):

And since you didn't mention it was perpendicular, I'm assuming that means it's parallel, so use the same slope.

OpenStudy (anonymous):

A?

OpenStudy (luigi0210):

No, no A

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

its not A? then is it D.?

OpenStudy (luigi0210):

Well you got good guessing skills :P

OpenStudy (luigi0210):

Yea, it's D.

OpenStudy (anonymous):

i always have. but im trying to understand i didnt completely guess lol

OpenStudy (luigi0210):

Well want an explanation?